# Violation of Quantum Bilocal Inequalities on Mutually-Commuting von Neumann Algebra Models

### 01 Apr 2026 v2

@@
Abstract.
In contrast to the non-relativistic quantum mechanics, the violation of Bell inequalities in quantum field theory depends more on the structure of observable algebras (typically type III von Neumann algebras) rather than the choice of specific quantum states. Therefore, studying the violation of Bell inequalities based on the von Neumann algebraic framework can be applied to addressing questions in quantum field theory, while also providing insights into the structural properties of von Neumann algebras. In this paper, we employ three mutually-commuting von Neumann algebras to characterize quantum entanglement swapping networks, and establish Bell-like inequalities thereon, commonly referred to as bilocal inequalities. We investigate the algebraic structural conditions under which bilocal inequalities are satisfied or violated on the generated algebra of these three von Neumann algebras. Furthermore, the conditions for maximal violation of the inequalities can be utilized to infer the structural information of von Neumann algebras in reverse.
Corresponding author
Key words and phrases: von Neumann algebra, Bell nonlocality, Bell inequality, Quantum network
1. Introduction
@@
In non-relativistic quantum mechanics, Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3]. Such nonlocal correlations can be detected through Bell inequalities, which serve as constraints that all local correlations must obey [4, 5, 6, 7, 8]. It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15]. However, it should be pointed out that relying solely on the non-relativistic quantum mechanics to discuss quantum information problems has two drawbacks [16, 17, 18]: 1) it fails to provide a universal framework for accurately describing phenomena in systems with infinite degrees of freedom; 2) it is not suitable for studying traditional quantum physics such as quantum field theory, which requires the language of type III von Neumann algebras. Research on quantum information problems on von Neumann algebras has received significant attention and yielded many meaningful results [19, 20, 21, 22, 23, 24, 25].
@@
In the 1980s early, mathematician Summers et al. introduce firstly the maximal violation of Bell inequalities and prove that its value is bounded by $2\sqrt${2} in the von Neumann algebra model, with equality attainable iff each subalgebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [26]. This shows that Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [27]. Translating these bounds into the vacuum representation of algebraic quantum field theory, they show that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges decay exponentially with mass-governed distance [28, 29, 30, 31, 32]. These works reveal a novel algebraic invariant, termed the Bell correlation invariant, which distinguishes infinitely many isomorphism classes of pairs of mutually commuting von Neumann algebras and links the maximal violation to the occurrence of the hyperfinite type $\rm II_{1}$ factor [27]. This is a pioneering work to make Bell nonlocality in quantum field theory (QFT) serve as a crucial bridge connecting quantum information science with fundamental physics [27, 33]. It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [34, 35, 36, 37, 38].
@@
In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states. Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [39, 40, 41, 42, 43, 44]. To date, Bell-like inequalities in the non-relativistic quantum mechanics have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [39, 45], chain configurations [46, 47], star topologies [48, 49], polygon structures [50, 51, 52], tree-shaped networks [53, 54, 55], arbitrary acyclic networks [41, 42, 56], and arbitrary $k$-independent networks [57]. Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [58, 59]. Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum network topologies [60]. Recently, the notion of bi-locality in an entanglement swapping network based on the mutually-commuting von Neumann algebra model has already been introduced by Ligthart et al. [61], and Xu has addressed the inclusion problem between the tensor product algebra model and the mutually-commuting von Neumann algebra model in this setting [62]. However, Bell-like inequalities for such von Neumann algebraic models have not yet been established. In this paper, we aim to establish bi-local inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
2. Mutually-commuting von Neumann algebra models, Tsirelson problem, and Connes’ embedding conjecture
In QFT, the algebras assigned to touching spacetime regions are of type III von Neumann ones, so there are no pure normal states and no tensor-product structures [63, 64, 65]. This differs from the non-relativistic quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [66, 67, 68]. These are two distinct models, referred to respectively as the tensor product algebra model and the mutually-commuting von Neumann algebra model.
In the tensor-product algebra model, a bipartite composite quantum system is described by the tensor product of two Hilbert spaces $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{B}}$, denoted as $\mathcal{H}=\mathcal{H}_{\mathcal{A}}\otimes\mathcal{H}_{\mathcal{B}}.$ Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all bounded linear operators on $\mathcal{H}$, and $\mathcal{T}(\mathcal{H})$ the space of trace-class operators. Then there is a duality relation $\mathcal{B}(\mathcal{H})=\mathcal{T}(\mathcal{H})^{*}$. A quantum state $\rho$ is a positive semidefinite operator with trace 1, and the set of states is denoted by $\mathcal{S}(\mathcal{H})=\{\rho\in\mathcal{T}(\mathcal{H})\mid\rho\geq 0,\ \mathrm{Tr}(\rho)=1\}$. The algebra of observables is the Jordan algebra of self-adjoint operators, denoted by $\mathcal{B}_{s}(\mathcal{H})=\{A\in\mathcal{B}(\mathcal{H})\mid A=A^{*}\}.$ In the mutually-commuting von Neumann algebra model, the algebra of observables of a quantum system is described by a von Neumann algebra $\mathcal{M}$, with $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ being two mutually commuting von Neumann subalgebras of $\mathcal{M}$ such that $(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}=\mathcal{M}$. Here, $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$.
However, in more general cases, a natural question arises: are the two models isomorphic? In other words, does $\mathcal{M}$ admit an isomorphism to $\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}$? This question is intimately related to the famous Tsirelson problem, which is equivalent to the Connes embedding conjecture in von Neumann algebra theory [69, 70, 71]. The Tsirelson problem asks whether the sets of joint probability distributions of measurement outcomes are the same in the two models. In 2020, Ji and collaborators [71] designed a quantum interactive protocol based on the classical halting problem in computability theory, and proved that the answer to the Tsirelson problem is negative. This means that the tensor-product algebra model and the mutually-commuting von Neumann algebra model are not equivalent. Therefore, the mutually-commuting von Neumann algebra model provides a more general framework, and many novel quantum phenomena fundamentally different from those in the tensor-product model have been discovered in this setting [71, 72, 73, 74].
@@ 2. Mutually-commuting von Neumann algebra models, Tsirelson problem, and Connes’ embedding conjecture
3. Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
@@
Figure 1. An entanglement swapping network scenario with two sources. The connection between two parties represents the sharing of the physical system between them.
@@
Quantum bilocal scenario. In non-relativistic quantum mechanics, the quantum entanglement swapping network (see Fig. 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them. The inputs and outputs of the measurements performed by the three parties are denoted as $x,y,z$ and $a,b,c$, respectively, where $x,y,z,a,b,c\in\{0,1\}$. Assume that each party performs binary-input and binary-output measurements, with the observables for Alice, Bob, and Charles denoted as $A_{x}$, $B_{y}$ and $C_{z}$, respectively. Here it is required that the spectra of operators $A_{x},B_{y},C_{Z}$ are all {-1,1}, implying that $I\leq A_{x},B_{y},C_{z}\leq I$. The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$. In this scenario, $p(abc|xyz)$ is said to be bilocal if it can be written as
@@
(3.1) $\mathcal{S}\equiv\sqrt{|I|}+\sqrt{|J|}\leq 2,$
@@
$\langle A_{x}B_{y}C_{z}\rangle=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}{\rm tr}((A_{a|x}B_{b|y}C_{x|z})\rho_{AB}\otimes\rho_{BC})=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz).$
Here $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$ and $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x},\ B_{b|y}$, and $C_{c|z}$ are the positive operator-valued measurement (POVMs) performed by Alice, Bob and Charles, respectively.
Furthermore, it follows from Ref. [57] that the violation of the inequality
(3.2) $\mathcal{S}^{\prime}=|K|+|L|\leq 2$
can also be used to detect nonbilocal correlations, where $K=\langle A_{0}(B_{0}-B_{1})C_{0}\rangle$ and $L=\langle A_{1}(B_{0}+B_{1})C_{1}\rangle$ [78]. Notably, different inequalities exhibit varying capabilities in detecting nonlocality.
@@
Definition 1. (Ternary Mutually-commuting von Neumann Algebra Models of Tripartite Quantum Systems) Let $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ be von Neumann subalgebras of $\mathcal{B}(\mathcal{H})$ over some Hilbert space $\mathcal{H}$, which are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$, where $\mathcal{M}_{j}^{\prime}$ is the communtant of $\mathcal{M}_{j}$. The generated von Neumann algebra
@@
We intend to use the above model of ternary mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig. 1. We note that there is no correlation between the parties $\mathcal{A}$ and $\mathcal{C}$ in the network. Mathematically, this independence can be described by the following formula. So the network state of the entanglement swapping network $\tau$ should be a state in the dual space $\mathcal{M}_{\mathcal{ABC}}^{*}$, satisfying
() $\tau(AC)=\tau(A)\tau(C)$
for any $A\in\mathcal{M}_{\mathcal{A}},C\in\mathcal{M}_{\mathcal{C}}$. We call it the independent condition (*). This assumption will be used throughout this paper.
Definition 2. A ternary mutually-commuting von Neumann algebra model of entanglement swapping networks is the ternary mutually-commuting von Neumann algebra model of tripartite quantum systems with the independence condition (*).
4. Bilocal inequalities and their bounds
In this section, we further analyze the conditions under which the bilocal inequality holds or is violated in the TMCvNA model of an entanglement swapping network. Specifically, in this model, we can construct bilocal inequalities analogous to those in the non-relativistic setting. Define
$I_{\tau}=\tau\bigl((A_{0}+A_{1})B_{0}(C_{0}+C_{1})\bigr),\ J_{\tau}=\tau\bigl((A_{0}-A_{1})B_{1}(C_{0}-C_{1})\bigr),$
where $\tau$ is the state on $\mathcal{M}_{\mathcal{ABC}}$ satisfying the independent condition (*). Here, $\tau\bigl(A_{x}B_{y}C_{z}\bigr)=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}\tau({A}_{a|x}{B}_{b|y}{C}_{c|z})$ and $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$, $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x}$, $B_{b|y}$, and $C_{c|z}$ are the POVMs performed by Alice, Bob, and Charles, respectively. Moreover, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ in the TMCvNA model is defined as

### 15 Apr 2026 v5

# B. Zheng, S. Yang, J. Hou and K. He

@@
[…]
B. Zheng, S. Yang, J. Hou and K. He
[…]
Violation of Bell-type Inequalities on Mutually-commuting von Neumann Algebra Models
College of Mathematics, Taiyuan University of Technology, Taiyuan, 030024, China
School of Mathematics, North University of China, Taiyuan, 030051, China,
School of Cyberspace Science and Technology, Beijing Institute of Technology, Beijing, 100081, China
E-mail :: yangshuyuan2000@163.com
College of Mathematics, Taiyuan University of Technology, Taiyuan, 030024, China
E-mail :: jinchuanhou@aliyun.com
College of Mathematics, Taiyuan University of Technology, Taiyuan, 030024, China
E-mail :: hekan@tyut.edu.cn
Violation of Bell-type Inequalities on Mutually-commuting von Neumann Algebra Models of Entanglement Swapping Networks  [^1] Shuyuan Yang and Bingke Zheng contribute equally to the work. Supported by the National Natural Science Foundation of China (Grant No. 12271394, 12571138)
Bingke ZHENG    Shuyuan YANG    Jinchuan HOU    Kan HE
[…]
Violation of Bell inequalities in bipartite systems represented by mutually-commuting von Neumann algebras has pioneered the study of vacuum entanglement in algebraic quantum field theory. It is unexpected that the maximal violation of Bell inequality can discover algebraic structures. In the paper, we establish the mutually-commuting von Neumann algebra model for entanglement swapping networks and Bell-type inequalities on this model. It generalizes the bipartite case to the ternary case. These algebras are all general von Neumann algebras, which provide a natural perspective to investigate Bell nonlocality in quantum networks in the infinitely-many-degree-of-freedom setting. We determine various bounds for Bell-type inequalities based on the structure of von Neumann algebras, and identify the algebraic structural conditions required for their violation. Finally, we show that the maximal violation of Bell-type inequalities in entanglement swapping networks can be used to determine partially the type classification of the underlying von Neumann algebras.
[…]
von Neumann algebra, Bell nonlocality, Bell inequality, Quantum network
[…]
81R15, 47A63, 47L30, 47N50
1 Introduction
@@
In non-relativistic quantum mechanics, Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3]. Such nonlocal correlations can be detected through Bell inequalities, which serve as constraints that all local correlations must obey [4, 5, 6, 7, 8]. It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15].
Meanwhile, motivated by the quantum field theory (QFT), which originates from the study of relativistic quantum mechanics, many novel quantum phenomena in systems with infinitely many degrees of freedom have been discovered [16, 17, 18, 19, 20, 21, 22]. This differs from the non-relativistic quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [23, 24, 25]. These are two distinct models, referred to respectively as the tensor product algebra (TPA) model and the mutually-commuting von Neumann (observable) algebra (MCvNA) model. In the MCvNA model, there is, in general, no tensor product decomposition of the Hilbert space describing subsystems. However, it should be pointed out that relying solely on the TPA model to discuss quantum information problems has drawbacks [26, 27, 28]. It fails to provide a universal framework for accurately describing phenomena in systems with infinite degrees of freedom and the quantum field theory, which requires the language of type III von Neumann algebras. Research on quantum information problems on von Neumann algebras has received significant attention and yielded many meaningful results from a mathematical perspective [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43].
@@
In the MCvNA model, the algebra of observables of quantum systems is described by a von Neumann algebra $\mathcal{M}$, with $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ being two mutually commuting von Neumann subalgebras of $\mathcal{M}$ such that $(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}=\mathcal{M}$. Here, $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$ [44, 45, 19]. It has been shown that the mutually-commuting von Neumann algebra model provides a more general framework [19]. In the 1980s early, Summers et al. first introduced the maximal violation of Bell inequality and proved that its value is bounded by $2\sqrt${2} in the MCvNA model of bipartite systems, with equality attainable if and only if each algebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [46]. This shows that Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [47]. Translating these bounds into the vacuum representation of algebraic quantum field theory, they show that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges decay exponentially with mass-governed distance [48, 49, 50, 51, 52]. These works reveal a novel algebraic invariant, termed the Bell correlation invariant, which distinguishes infinitely many isomorphism classes of pairs of mutually commuting von Neumann algebras and links the maximal violation to the occurrence of the hyperfinite type $\rm II_{1}$ factor [47]. This is a pioneering work to make Bell nonlocality in QFT serve as a crucial bridge connecting quantum information science with fundamental physics [47, 53]. It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [54, 55, 56, 57, 58].
@@
In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states. Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [59, 60, 61, 62, 63, 64]. To date, Bell-type inequalities in the non-relativistic quantum mechanics have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [59, 65], chain configurations [66, 67], star topologies [68, 69], polygon structures [70, 71, 72], tree-shaped networks [73, 74, 75], arbitrary acyclic networks [61, 62, 76], and arbitrary $k$-independent networks [77]. Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [78, 79]. Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum network topologies [80]. Recently, the notion of bilocality in an entanglement swapping network based on the MCvNA model has already been introduced by Ligthart et al. [81, 82], and Xu has addressed the inclusion problem between TPA model and MCvNA model in this setting [83]. However, Bell-type inequalities in the MCvNA model have not yet been established. In this paper, we aim to establish bilocal inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
@@
2 Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
@@
Figure 1: An entanglement swapping network scenario with two sources. The connection between two parties represents the sharing of the physical system between them.
@@
Quantum bilocal scenario. In non-relativistic quantum mechanics, the quantum entanglement swapping network (see Fig. 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them. The inputs and outputs of the measurements performed by the three parties are denoted as $x,y,z$ and $a,b,c$, respectively. Assume that each party performs binary-input and binary-output measurements, with the observables for Alice, Bob, and Charles denoted as $A_{x}$, $B_{y}$ and $C_{z}$, respectively, with $x,y,z,a,b,c\in\{0,1\}$. Here it is required that the spectra of operators $A_{x},B_{y},C_{z}$ are all $\{-1,1\}$, implying that $I\leq A_{x},B_{y},C_{z}\leq I$. The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$. In this scenario, $p(abc|xyz)$ is said to be bilocal if it can be written as
@@
$\mathcal{S}\equiv\sqrt{|I|}+\sqrt{|J|}\leq 2,$ (2.1)
@@
$\langle A_{x}B_{y}C_{z}\rangle$ $=$ $\sum_{a,b,c=0}^{1}(-1)^{a+b+c}{\rm tr}((A_{a|x}B_{b|y}C_{x|z})\rho_{AB}\otimes\rho_{BC})$
$=$ $\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz).$
Here $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$ and $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x},\ B_{b|y}$, and $C_{c|z}$ are the positive operator-valued measurements (POVMs) performed by Alice, Bob and Charles, respectively.
@@
Definition 2.1.
(Ternary Mutually-commuting von Neumann Algebra Models of Tripartite Quantum Systems) Let $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ be von Neumann subalgebras of $\mathcal{B}(\mathcal{H})$ over some Hilbert space $\mathcal{H}$, which are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$, where $\mathcal{M}_{j}^{\prime}$ is the communtant of $\mathcal{M}_{j}$. The generated von Neumann algebra
@@
We intend to use the above model of ternary mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig. 1. We note that there is no correlation between the parties Alice and Charles in the network. Mathematically, this independence can be described by the following formula: a network state $\tau$ of the entanglement swapping network should be a state in $\mathcal{M}_{\mathcal{ABC}}^{*}$, the dual space of $\mathcal{M}_{\mathcal{ABC}}$, satisfying
$\tau(AC)=\tau(A)\tau(C)$ ()
for any $A\in\mathcal{M}_{\mathcal{A}},C\in\mathcal{M}_{\mathcal{C}}$. We call (*) the independent condition. This assumption will be used throughout this paper.
Definition 2.2.
The ternary mutually-commuting von Neumann algebra model of entanglement swapping networks is the ternary mutually-commuting von Neumann algebra model of tripartite quantum systems with all states satisfying the independence condition (*).
3 Bilocal inequalities and their bounds
In this section, we further analyze the conditions under which the bilocal inequality holds or is violated in the TMCvNA model of an entanglement swapping network. Specifically, in this model, we can construct the bilocal inequality analogous to that in the non-relativistic setting. Let
$I_{\tau}=\tau\bigl((A_{0}+A_{1})B_{0}(C_{0}+C_{1})\bigr),$
$J_{\tau}=\tau\bigl((A_{0}-A_{1})B_{1}(C_{0}-C_{1})\bigr),$
where $\tau$ is a state on $\mathcal{M}_{\mathcal{ABC}}$ satisfying the independent condition (*). Here, $\tau\bigl(A_{x}B_{y}C_{z}\bigr)=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}\tau({A}_{a|x}{B}_{b|y}{C}_{c|z})$ and $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$, $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x}$, $B_{b|y}$, and $C_{c|z}$ are the POVMs performed by Alice, Bob, and Charles, respectively. Moreover, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ in the TMCvNA model is defined as